# Breaking through the learning plateaus of in-context learning in Transformer

### v2

@@ E.4 Performance analysis
The probability for a in-context example having same value of a factor as prompt sample is $\frac{1}{n_{v}}$ and the probability of having different values is $\frac{n_{v}-1}{n_{v}}$. Therefore, given $i$ samples, the probability for $j$ samples have a same value of a factor as prompt sample is $\tbinom{i}{j}\frac{(n_{v}-1)^{i-j}}{n_{v}^{i}}$.
@@
Given $i$ in-context examples, the probability for we cannot distinguish $k$ factors is
@@
When we cannot distinguish the hidden factor from $k$ factors, the probability to predict wrong results is $\frac{k-1}{k}$. Combining the results above, we obtain the error that
@@
The probability to give a right prediction is
@@
In the constructed Transformer, we will autogressively combining the results of the previous prediction (Corresponding to Layer 2), we have:
@@
The copy-past mechanism is used to predict the answer of the prompt example (Corresponding to layer 3). For the copy-past mechanism, have a in-context example with same prediction result as the prompt example is neccesary. When we correct predict the hidden factor, the probility to predict correctly is $1-(\frac{n_{v}-1}{n_{v}})^{i}$. When we predict a wrong hidden factor, the probility is $\frac{1}{n_{v}}$. Combine the two above, we obtain the accuracy

### v3

@@ E.4 Performance analysis
The probability for an in-context example having the same value of a factor as the prompt sample is $\frac{1}{n_{v}}$ and the probability of having different values is $\frac{n_{v}-1}{n_{v}}$. Therefore, given $i$ samples, the probability for $j$ samples have the same value of a factor as the prompt sample is $\tbinom{i}{j}\frac{(n_{v}-1)^{i-j}}{n_{v}^{i}}$.
@@
Given $i$ in-context examples, the probability for that we cannot distinguish $k$ factors is
@@
When we cannot distinguish the hidden factor from $k$ factors, the probability of predicting wrong results is $\frac{k-1}{k}$. Combining the results above, we obtain the error:
@@
The probability of giving the right prediction is
@@
In the constructed Transformer, we will autoregressively combine the results of the previous prediction (Corresponding to Layer 2). We have:
@@
The copy-past mechanism is used to predict the answer to the prompt example (Corresponding to layer 3). For the copy-past mechanism, having an in-context example with the same prediction result as the prompt example is necessary. When we correctly predict the hidden factor, the probability to predict correctly is $1-(\frac{n_{v}-1}{n_{v}})^{i}$. When we predict a wrong hidden factor, the probability is $\frac{1}{n_{v}}$. Combine the two above, we obtain the accuracy as follows